# Ermakov–Lewis Invariants in Stationary Bohm–Madelung Quantum Mechanics

### 09 Mar 2026 v2

@@ 1 Introduction
The stationary Bohm–Madelung(BM) formulation is discussed in Holland’s Quantum Theory of Motion [1], where the one-dimensional amplitude equation is noted to be formally related to an Ermakov-type differential equation. This observation, however, is not developed into a general invariant framework. In the present work, we show that Ermakov–Pinney (EP) equations and their invariants arise naturally and systematically in stationary quantum mechanics for diagonal, separable Hamiltonians.
@@
The present work addresses this gap by demonstrating that invariants arise naturally in stationary quantum mechanics whenever the Hamiltonian is diagonal and separable, with the Ermakov–Pinney (EP) equation emerging as the amplitude-space counterpart of Hamilton–Jacobi (HJ) separability. In this setting, the spatial coordinate plays the role of an evolution parameter, and invariant structure appears at the level of stationary amplitude dynamics rather than through explicit time dependence.
@@
Here, we show that this distinction is largely historical rather than structural. When stationary quantum systems are expressed in the BM representation, the same mathematical structure underlying EL invariants emerges intrinsically from the stationary continuity constraint and the separability of the Hamiltonian. For diagonal and separable Hamiltonians, this structure leads systematically to EP equations governing the amplitude dynamics in each degree of freedom, together with conserved EL invariants.

### 19 Mar 2026 v3

@@ 1 Introduction
The stationary Bohm–Madelung (BM) formulation is discussed in Holland’s The Quantum Theory of Motion [1], where the one-dimensional amplitude equation is noted to be formally related to an Ermakov-type differential equation. This observation, however, is not developed into a general invariant framework. In the present work, we show that Ermakov–Pinney (EP) equations and their invariants arise naturally and systematically in stationary quantum mechanics for diagonal, separable Hamiltonians.
@@
The present work addresses this gap by demonstrating that invariant structures arise naturally in stationary quantum mechanics whenever the Hamiltonian is diagonal and separable, with the Ermakov–Pinney (EP) equation emerging as the amplitude-space counterpart of Hamilton–Jacobi (HJ) separability. In this setting, the spatial coordinate plays the role of an evolution parameter, and the invariant structure appears at the level of stationary amplitude dynamics rather than through explicit time dependence.
@@
Ermakov–Lewis invariant structures in quantum mechanics have been discussed in earlier works, particularly by Reinisch [12] and Schuch [13, 14] in Bohm–Madelung and Riccati formulations. The present work extends the formalism to separable coordinate systems and demonstrates its application by analytically solving the Born–Oppenheimer two-center potential.
We show that this distinction is largely historical rather than structural. When stationary quantum systems are expressed in the BM representation, the same mathematical structure underlying EL invariants emerges intrinsically from the stationary continuity constraint and the separability of the Hamiltonian. For diagonal and separable Hamiltonians, this structure leads systematically to EP equations governing the amplitude dynamics in each degree of freedom, together with conserved EL invariants.
@@
Given the broad scope of the present work, spanning variational formulations, coordinate-separable structures, and nonlinear representations of quantum mechanics, it is possible that some related contributions have not been cited. Any such omissions are unintentional, and the authors welcome further references to relevant literature.